# CoD: A Diffusion Foundation Model for Image Compression

### v1

@@ Appendix E Additional Details and Results
E.1 Training Details
CoD Low-Resolution Pre-Training $256\times 256$ 0.0156 bpp / 1024 bits 22.1M 400 K $4\times 32$ $1\times 10^{-4}$ 67
CoD High-Resolution Pre-Training $512\times 512$ 0.0039 bpp / 1024 bits 12.8 M 100 K $4\times 16$ $2\times 10^{-5}$ 25
CoD Unified Post-Training $512\times 512$ 0.0039 bpp / 1024 bits 12.8 M 50 K $4\times 16$ $2\times 10^{-5}$ 24
Table 3: Detailed configuration of each CoD training stage. The full training process takes 116 A100 GPU hours (approximately 5 days).
E.1.1 Datasets
@@ E.1.2 Unified Training
E.1.3 Multi-Stage Training
@@
E.2 Complexity Analysis
Table 5 compares the computational cost of Stable Diffusion v1.5 and CoD. Stable Diffusion requires a large captioning model to generate text conditions, whereas CoD relies on a lightweight 177M image encoder and decoder, taking only 24 ms to process a $512\times 512$ image. The diffusion module in CoD also incurs lower latency and parameter overhead. Pixel-space CoD is slightly slower than latent-space CoD due to the additional decoupled pixel head, but it avoids the expensive VAE decoding, which makes it faster in the few-step regime.
Despite being faster than Stable Diffusion, the inference speed of CoD remains slow to meet the needs of real-time coding. This limitation can be mitigated by distilling the diffusion process into fewer steps and smaller models. Section B shows the potential of one-step CoD, and Section C demonstrates that CoD can be used as a perceptual supervisory signal. Together, these findings suggest a promising path toward fast inference: train a lightweight CoD and distill it into a single step using the large CoD as perceptual supervision. We leave this direction to future work with the goal of real-time diffusion codecs.
Speed (ms) / Params. Conditioner Diffusion Module
Diffusion VAE Decoder Total (25 steps)
Stable Diffusion v1.5 203.0 / 3.7 B* 30.6 / 860 M 43.4 / 49 M 808.4 / 909 M
Latent-space CoD 24.0 / 177 M 21.5 / 676 M 43.4 / 49 M 580.9 / 724 M
Pixel-space CoD 24.0 / 177 M 25.5 / 720 M 0 / 0 M 637.5 / 720 M
E.3 Visual Comparison

### v2

@@ Appendix F Additional Details and Results
F.1 Training Details
F.1.1 Datasets
@@ F.1.2 Unified Training
CoD Low-Resolution Pre-Training $256\times 256$ 0.0156 bpp / 1024 bits 22.1M 400 K $4\times 32$ $1\times 10^{-4}$ $4\times 67$
CoD High-Resolution Pre-Training $512\times 512$ 0.0039 bpp / 1024 bits 12.8 M 100 K $4\times 16$ $2\times 10^{-5}$ $4\times 25$
CoD Unified Post-Training $512\times 512$ 0.0039 bpp / 1024 bits 12.8 M 50 K $4\times 16$ $2\times 10^{-5}$ $4\times 24$
Table 4: Detailed configuration of each CoD training stage. The full training process takes 464 A100 GPU hours (approximately 20 days).
Speed (ms) / Params. Per-Module Breakdown Steps Total
Conditioner Diffusion VAE Decoder
Stable Diffusion v1.5 203.0 / 3.7 B* 30.6 / 860 M 43.4 / 49 M 25 1011 / 4.6 B
Latent-space CoD 8.3 / 177 M 21.5 / 676 M 43.4 / 49 M 25 589.2 / 901 M
Pixel-space CoD 8.3 / 177 M 25.5 / 720 M - / - 25 645.8 / 897 M
One-Step Pixel-space CoD 8.3 / 177 M 16.9 / 513 M** - / - 1 25.2 / 690 M
    The AdaLN-Zero layers are omitted since they become redundant when inference operates with a fixed $t=0$.
F.1.3 Multi-Stage Training
@@
F.2 Complexity Analysis
Table 5 compares the computational cost of Stable Diffusion v1.5 and CoD. Stable Diffusion requires a large captioning model to generate text conditions, whereas CoD relies on a lightweight 177M image encoder and decoder, taking only 8.3 ms to process a $512\times 512$ image. The diffusion module in CoD also incurs lower latency and parameter overhead. Pixel-space CoD is slightly slower than latent-space CoD due to the additional decoupled pixel head, but it avoids the expensive VAE decoding, which makes it faster in the few-step regime.
Compared to multi-step diffusion codecs, the one-step pixel-space CoD offers significantly faster inference with a much smaller parameter footprint (see Section B). It enables real-time coding, achieving $25.2$ ms per $512\times 512$ image. In the future, we hope it can be further accelerated by training a smaller model. Section C demonstrates that CoD can be used as a perceptual supervisory signal, suggesting a promising path toward fast inference: train a lightweight CoD and distill it into a single step using the large CoD as perceptual supervision. We leave this direction to future work with the goal of real-time high-resolution diffusion codecs.
F.3 Visual Comparison
@@
Figure 18: More visualization results on CLIC 2020 test set [clic].